OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

let Tim's age be x , then Johnson age will be x+10 .... When Tim started football 3 years ago his age was x-3 which equals (x+10-3)/2... solve the equation

OpenStudy (anonymous):

thankyou , what about ......... Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (phi):

Video on solving equations. The following two are also helpful http://www.khanacademy.org/math/algebra/solving-linear-equations/v/simple-equations

OpenStudy (phi):

x - 62,500 = 0.25y y is the total number of votes=125,000 people voted replace y in the equation with its value x- 62500= 0.25*125000 to solve, add 62500 to both sides of the equation (watch the video if this isn't totally clear) x= 0.25*125000+62500 now it is arithmetic, multiply and add.